tmmicheal:

Binh solved this system of equations by graphing. Line 1: y = negative one-half x minus three-halves. Line 2: y = x minus 3. Binh’s Graph On a coordinate plane, a line with equation y = x minus 3 goes through (negative 3, 0) and (0, 3). A line with equation y = negative one-half x minus three-halves goes through (negative 3, 0) and (1, negative 2). Binh says the point of intersection is (0, –3). Which statements identify the errors Binh made? Check all that apply. Binh listed the coordinates in the wrong order when describing the point of intersection on his graph. Binh should have graphed the y-intercept of y = x minus 3 at (0, negative 3). Binh should have graphed the y-intercept of y = x minus 3 at (0, 1). Binh should have graphed the y-intercept of y = negative one-half x minus three-halves at (0, negative one-half). Binh should have found the point of intersection to be (1, negative 2)

Shadow:

\[y_{1} = -\frac{ 1 }{ 2 }x - \frac{ 3 }{ 2 }\] \[y_{2} = x - 3\]

Shadow:

@tmmicheal Have you graphed them out?